OpenStudy (anonymous):

What is the measure of the vertex angle of an isosceles triangle if one of its base angles measures 42°? 96° 84° 69° 138°

OpenStudy (anonymous):

Its 90 degrees coz the sum of all angles is 180 and since base angles are equal we hav 2 angles of total 42+42=84 degrees hence vertex angle = 180 - 84 = 96 credit: found it on yahoo answers https://answers.yahoo.com/question/index?qid=20080122072756AAkBdqC